# Training Versatile Coding Agents in Synthetic Environments

### v1

@@ 1 Introduction
Recent advances in Large Language Models (LLMs) have empowered Language Model Agents (LM Agents) to address a wide array of real-world tasks, with coding being a prime example (jimenez2024swebench; zhuo2024bigcodebench; jain2024livecodebench). In an effort to enhance the coding proficiency of LLMs and agents, prior studies (pan2025training; jain2025r2e; yang2025swesmith) have explored creating specialized training environments designed to develop state-of-the-art open-weight software engineering (SWE) agents. These approaches utilize automated pipelines to collect SWE tasks with verifiable reward signals from existing Github repositories, enabling the training of SWE agents through agentic distillation or reinforcement learning (RL) (deepswe2025).
However, as pan2025training has demonstrated, the performance of SWE agents is highly dependent on the scale of the training data. Though pipelines of jain2025r2e and yang2025swesmith mitigate the reliance on human-written pull requests, they remain fundamentally constrained by their reliance on existing GitHub repositories, thereby hindering the potential of creating larger-scale training datasets while maintaining high quality and flexibility. Furthermore, real-world coding spans a diverse spectrum of tasks, including resolving issues (jimenez2024swebench), generating unit tests to detect or reproduce potential problems (mundler2024swtbench), and developing new projects from scratch (zhao2024commit0librarygenerationscratch). An ideal coding agent should be capable of handling these varied tasks, yet all aforementioned works predominantly focus on issue-resolution tasks while neglecting other critical scenarios. As illustrated in Figure 1, agents trained in these environments exhibit less pronounced or even deteriorated performance on other well-established benchmarks, despite their impressive results on SWE-bench.
@@
SWE-Playground supports training versatile coding agents. Using trajectories collected exclusively from SWE-Playground, we train models to validate the effectiveness of our approach. Our model consistently demonstrates superior performance over the base model across three distinct benchmarks that assess a wide range of coding capabilities. In stark contrast, agents trained on prior environments, such as SWE-Gym (pan2025training) and R2E-Gym (jain2025r2e), which are typically limited to SWE-bench format tasks, exhibit limited generalization to these benchmarks. While they show minor gains on some out-of-domain tasks, these improvements do not match their source domain performance, and they suffer noticeable degradation on others. This result highlights that though prior environments successfully produce specialized agents that excels at SWE-bench, our SWE-Playground is capable of cultivating robust and generalizable coding proficiency, and thus developing versatile coding agents.
SWE-Playground produces trajectories with dense training signal. Our experiments collect 704 trajectories, a set significantly smaller than those used by R2E-Gym (jain2025r2e) and SWE-smith (yang2025swesmith). Despite this smaller dataset, our models achieve comparable performance. Further analysis reveals that our trajectories are substantially richer, containing more tokens and tool calls on average compared to those from previous environments. Moreover, they feature a higher proportion of bash execution, demonstrating a paradigm of execution-based software development rather than simple code generation. We attribute this high data efficiency to the dense and complex nature of SWE-Playground tasks and generated trajectories.
@@ 2.1 Necessary Skills and Evaluation Benchmarks
Software engineering is a multifaceted discipline that extends far beyond merely writing code, encompassing tasks such as designing, coding, testing, and debugging. Underscoring this, meyer2019today studies the work of human software developers and finds that only 15% of their time is spent actively coding. Moreover, code development itself is not monolithic, and it requires a diverse set of specialized skills for effective software engineering.
Fortunately, there are a variety of benchmarks designed to evaluate the capabilities of language models to tackle these various coding tasks. SWE-bench (jimenez2024swebench) is the canonical benchmark for the “issue in, patch out” bug-fixing use case, focusing on popular Python repositories. SWE-bench-Live (zhang2025swebenchgoeslive) addresses the static limitations of SWE-bench, introducing a live and automated pipeline that curates fresh, contamination-resistant tasks from real-world GitHub issues. SWE-Bench Pro (deng2025swebenchproaiagents) also builds upon SWE-bench with more challenging tasks to capture realistic, complex, enterprise-level problems beyond the scope of SWE-bench.
@@
  • Building Libraries from Scratch: Commit-0 (zhao2024commit0librarygenerationscratch) tasks models with rebuilding and generating the entire library from the initial commit.
  • Multilingual Bug Fixing: Multi-SWE-bench  (zan2025multiswebench) and SWE-bench Multilingual (yang2025swesmith) cover a wider range programming languages beyond Python, including C++, TypeScript, Rust, etc.
  • Multimodal Understanding: ArtifactsBench (zhang2025artifactsbenchbridgingvisualinteractivegap) and SWE-bench Multimodal (yang2024swebenchmultimodal) test the capability to ground code generation and understanding in visual artifacts.
  • Test Generation: SWT-Bench (mundler2024swtbench) and TestGenEval (jain2025testgenevalrealworldunit) evaluates model performance on reproducing GitHub issues in Python.
  • Performance Optimization: SWE-Perf (he2025sweperf), SWE-fficiency (ma2025swefficiencylanguagemodelsoptimize), and KernelBench (ouyang2025kernelbenchllmswriteefficient) benchmark models on optimizing code performance in diverse scenarios and programming languages.
@@ 2.2 Training Environments
Several recent works have focused on developing specialized environments for training SWE agents. SWE-Gym (pan2025training) first explores the construction of environments for training SWE agents with pre-installed dependencies and executable test verification derived from real-world GitHub issues. R2E-Gym (jain2025r2e) proposes a synthetic data curation recipe to create execution environments from commits through backtranslation and test collection or generation, which removes the reliance on human-written pull requests or unit tests. Similarly, SWE-smith (yang2025swesmith) facilitates automatic generation of scalable training environments and instances via function rewriting and bug combination. SWE-Mirror (wang2025swemirrorscalingissueresolvingdatasets) further scales this by mirroring issues across repositories. SWE-Flow (zhang2025sweflow) utilizes Test-Driven Development (TDD) to produce structured development tasks with explicit goals directly from unit tests, thereby generating training instances.
@@ 3.2 Task Proposal
Once a task is defined, we generate a corresponding detailed checklist that provides a granular explanation and explicitly specifies the requisite unit tests, standard cases, specific assertions, and potential edge cases. This practice offers two primary advantages. First, unit tests serve as a reliable reward signal. Though previous works have explored training software engineering verifier models for inference-time scaling (pan2025training; jain2025r2e), unit tests remain more trustworthy provided that current language models and agents can generate adequate tests (chou2025autocodebenchlargelanguagemodels). Second, the detailed checklist acts as a clear guide, steering the agent in implementing unit tests in the later stage.
@@ 3.5 Functionality Implementation
After the agent completes the implementation, the test suites in the repository are replaced with the original tests produced during the test generation stage, which is a critical step to mitigate the risk of reward hacking, where the agent might otherwise pass tests by directly editing the test code. The original unit tests are subsequently executed to verify the correctness of implementation. This overall process leverages a similar idea to the concept of self-questioning (chen2025selfquestioninglanguagemodels), as unit test generation and functionality implementation form a mutual verification loop. An issue in either stage, be it a flawed test or an incorrect implementation, will lead to the failure of unit tests, thereby adding to trustworthiness of the generated unit tests. and also robustness of the entire verification process. Notably, we alternate between unit test generation and functionality implementation, forming a loop. This approach ensures that subsequent unit tests are adapted to the the cumulative functionality, which simplifies the implementation the implementation process by removing the need of revisiting or modifying the previously validated code.
@@
Qwen2.5-Coder-7B - 01.81.8   +0.00+0.00 $45.8$   +0.00+0.00 00.720.72   +0.00+0.00 07.557.55   +00.00+00.00 05.215.21   +0.00+0.00
SWE-Gym-7B 491 $10.6$   $\phantom{0}8.8$ $33.8$   12.0-12.0 01.451.45   +0.73+0.73 05.195.19   $\phantom{0}2.36$ 07.197.19   +1.98+1.98
R2E-Gym-7B 3.3k $19.0$   +17.2+17.2 00.720.72 +   $0.00$ 02.662.66   $\phantom{0}4.89$ 07.287.28   +2.07+2.07
SWE-smith-7B 2.0k $15.2$   +13.4+13.4 00.000.00   0.72-0.72 $12.30$   $\phantom{0}4.75$ 07.507.50   +2.29+2.29
SWE-Play-mix-7B 704 $17.0$   +15.2+15.2 06.46.4   39.4-39.4 04.484.48   +3.76+3.76 $24.15$   +16.60+16.60 08.468.46   +3.25+3.25
@@
Qwen2.5-Coder-32B - 07.07.0   +0.00+0.00 09.59.5   +0.00+0.00 09.429.42   +0.00+0.00 $24.81$   +00.00+00.00 06.626.62   +0.00+0.00
SWE-Gym-32B 491 $20.6$   +13.6+13.6 $13.8$   $\phantom{0}4.3$ 03.263.26   6.16-6.16 $10.34$   14.47-14.47 07.197.19   +0.57+0.57
R2E-Gym-32B 3.3k $34.4$   +27.4+27.4 03.263.26   6.16-6.16 $15.04$   $\phantom{0}9.77$ 09.459.45   +2.83+2.83
SWE-smith-32B 5.0k $40.2$   +33.2+33.2 $13.77$   +4.35+4.35 $42.15$   +17.34+17.34 07.507.50   +0.88+0.88
SWE-Play-mix-32B 704 $31.2$   +24.2+24.2 06.86.8   $\phantom{0}2.7$ $18.12$   +8.70+8.70 $44.24$   +19.43+19.43 $10.42$   +3.80+3.80
@@ 4 Capability-Specific Adaption
Issue Resolution: SWE-bench (jimenez2024swebench) is the canonical benchmark for software engineering tasks, and it features resolving real GitHub issues. This task setting requires a repository that is generally functional but contains a minor bug affecting certain functionalities. With our pipeline, a functional repository already exists, with the only remaining task being the generation of a specific issue and its injection into the code. Similar to yang2025swesmith, we first leverage an LLM to formulate an issue based on the unit test documentation and code of the task, and then call an agent to inject the issue into the repository. The existing test suites can directly verify the correctness of the injection, where a failing test serves as the success signal.
Test Generation: SWT-Bench (mundler2024swtbench) is a test generation benchmark where agents are required to develop unit tests to reproduce issues. For this setting, beyond injecting the issue into the repository, we also require the agent to modify the test suites so that the injected bug is no longer directly detectable. The agents are then tasked with formulating a test script that exposes the faulty behavior and then integrate this script into the existing test suites to complete the task, a procedure identical to SWT-Bench.
Generating Libraries from Scratch: Commit-0 (zhao2024commit0librarygenerationscratch) evaluates agents on constructing the entire repository from scratch, with no step-to-step decomposition. For this setting, the complete and functional repositories generated by our pipeline serve as the target artifacts and reference implementation. To simulate the Commit-0 task format, we replace all function bodies with pass statements and then prompt agents to re-implement the project. Notably, we do not use the existing test suites to verify implementation correctness in our experiments. Instead, we let the agent attempt the full task and directly collect its trajectory, which aligns with a pure distillation setting. We adopt such strategy out of the practical necessity. Because our pipeline separates functionalities into step-by-step tasks and the unit tests are designed to be inherently stringent, it is challenging for the agent to pass all tests within a single call.
@@ 5.1 Experimental Settings
With our proposed SWE-Playground pipeline, we collect from 28 generated projects a total of 280 general trajectories (as detailed in Section 3.1 to Section 3.5) and 424 task specific trajectories (as detailed in Section 4), respectively 213 for issue resolution, 183 for issue reproduction and 28 for library generation. Following pan2025training and jain2025r2e, we use OpenHands (wang2025openhands) CodeAct Agent our agent Scaffold, and select Qwen-2.5-Coder series (hui2024qwen2) as our the base model. We adopt SWE-bench Verified (chowdhury2024swebenchverified), SWT-Bench Lite and Commit-0 Lite as the evaluation splits for the respective benchmarks. We provide details in Appendix B.
@@
Qwen2.5-Coder 01.81.8 $45.8$ $0.72$ 07.557.55 $5.21$
SWE-Play-mix $17.0$ 06.46.4 $4.48$ $24.15$ $8.46$
SWE-Play-general 08.68.6 $22.6$ $2.54$ 07.557.55 $7.50$
SWE-Play-swe $11.0$ $10.6$ $0.72$ 04.804.80 $6.81$
@@ 6.2 Investigation into Data Efficiency
These outcomes raise a critical question: why can our framework demonstrate better data efficiency compared with various previous methods? To answer this, we perform a detailed comparative analysis of our trajectories against those from baseline studies, with the results presented in Table 4. Our analysis reveals that trajectories generated with SWE-Playground are quantifiably more complex and comprehensive. As shown in the table, they contain two to three times more tokens and messages on average, including a higher number of assistant tokens, tool calls and lines of code edited, compared to other datasets. We characterize this quality as a higher learning density. This concept, in line with the Agency Efficiency Principle introduced by xiao2025limiagency, suggests that trajectories dense with diverse and complex actions provide a more potent learning impetus per sample. The richness of our trajectories demonstrates this principle in practice, leading to greater data efficiency.

### v2

@@ 1 Introduction
Recent advances in Large Language Models (LLMs) have empowered Language Model Agents (LM Agents) to address a wide array of real-world tasks, with coding being a prime example (Jimenez et al., 2024; Zhuo et al., 2024; Jain et al., 2024). In an effort to enhance the coding proficiency of LLMs and agents, prior studies (Pan et al., 2025; Jain et al., 2025b; Yang et al., 2025b) have explored creating specialized training environments designed to develop state-of-the-art open-weight software engineering (SWE) agents. These approaches utilize automated pipelines to collect SWE tasks with verifiable reward signals from existing Github repositories, enabling the training of SWE agents through agentic distillation or reinforcement learning (RL) (Luo et al., 2025).
However, as Pan et al. (2025) has demonstrated, the performance of SWE agents is highly dependent on the scale of the training data. Though pipelines of Jain et al. (2025b) and Yang et al. (2025b) mitigate the reliance on human-written pull requests, they remain fundamentally constrained by their reliance on existing GitHub repositories, thereby hindering the potential of creating larger-scale training datasets while maintaining high quality and flexibility. Furthermore, real-world coding spans a diverse spectrum of tasks, including resolving issues (Jimenez et al., 2024), generating unit tests to detect or reproduce potential problems (Mündler et al., 2024), and developing new projects from scratch (Zhao et al., 2024). An ideal coding agent should be capable of handling these varied tasks, yet all aforementioned works predominantly focus on issue-resolution tasks while neglecting other critical scenarios. As illustrated in Figure 1, agents trained in these environments exhibit less pronounced or even deteriorated performance on other well-established benchmarks, despite their impressive results on SWE-bench.
@@
SWE-Playground supports training versatile coding agents. Using trajectories collected exclusively from SWE-Playground, we train models to validate the effectiveness of our approach. Our model consistently demonstrates superior performance over the base model across three distinct benchmarks that assess a wide range of coding capabilities. In stark contrast, agents trained on prior environments, such as SWE-Gym (Pan et al., 2025) and R2E-Gym (Jain et al., 2025b), which are typically limited to SWE-bench format tasks, exhibit limited generalization to these benchmarks. While they show minor gains on some out-of-domain tasks, these improvements do not match their source domain performance, and they suffer noticeable degradation on others. This result highlights that though prior environments successfully produce specialized agents that excels at SWE-bench, our SWE-Playground is capable of cultivating robust and generalizable coding proficiency, and thus developing versatile coding agents.
SWE-Playground produces trajectories with dense training signal. Our experiments collect 704 trajectories, a set significantly smaller than those used by R2E-Gym (Jain et al., 2025b) and SWE-smith (Yang et al., 2025b). Despite this smaller dataset, our models achieve comparable performance. Further analysis reveals that our trajectories are substantially richer, containing more tokens and tool calls on average compared to those from previous environments. Moreover, they feature a higher proportion of bash execution, demonstrating a paradigm of execution-based software development rather than simple code generation. We attribute this high data efficiency to the dense and complex nature of SWE-Playground tasks and generated trajectories.
@@ 2.1 Necessary Skills and Evaluation Benchmarks
Software engineering is a multifaceted discipline that extends far beyond merely writing code, encompassing tasks such as designing, coding, testing, and debugging. Underscoring this, Meyer et al. (2019) studies the work of human software developers and finds that only 15% of their time is spent actively coding. Moreover, code development itself is not monolithic, and it requires a diverse set of specialized skills for effective software engineering.
Fortunately, there are a variety of benchmarks designed to evaluate the capabilities of language models to tackle these various coding tasks. SWE-bench (Jimenez et al., 2024) is the canonical benchmark for the “issue in, patch out” bug-fixing use case, focusing on popular Python repositories. SWE-bench-Live (Zhang et al., 2025c) addresses the static limitations of SWE-bench, introducing a live and automated pipeline that curates fresh, contamination-resistant tasks from real-world GitHub issues. SWE-Bench Pro (Deng et al., 2025) also builds upon SWE-bench with more challenging tasks to capture realistic, complex, enterprise-level problems beyond the scope of SWE-bench.
@@
  • Building Libraries from Scratch: Commit-0 (Zhao et al., 2024) tasks models with rebuilding and generating the entire library from the initial commit.
  • Multilingual Bug Fixing: Multi-SWE-bench  (Zan et al., 2025) and SWE-bench Multilingual (Yang et al., 2025b) cover a wider range programming languages beyond Python, including C++, TypeScript, Rust, etc.
  • Multimodal Understanding: ArtifactsBench (Zhang et al., 2025a) and SWE-bench Multimodal (Yang et al., 2025a) test the capability to ground code generation and understanding in visual artifacts.
  • Test Generation: SWT-Bench (Mündler et al., 2024) and TestGenEval (Jain et al., 2025a) evaluates model performance on reproducing GitHub issues in Python.
  • Performance Optimization: SWE-Perf (He et al., 2025), SWE-fficiency (Ma et al., 2025), and KernelBench (Ouyang et al., 2025) benchmark models on optimizing code performance in diverse scenarios and programming languages.
@@ 2.2 Training Environments
Several recent works have focused on developing specialized environments for training SWE agents. SWE-Gym (Pan et al., 2025) first explores the construction of environments for training SWE agents with pre-installed dependencies and executable test verification derived from real-world GitHub issues. R2E-Gym (Jain et al., 2025b) proposes a synthetic data curation recipe to create execution environments from commits through backtranslation and test collection or generation, which removes the reliance on human-written pull requests or unit tests. Similarly, SWE-smith (Yang et al., 2025b) facilitates automatic generation of scalable training environments and instances via function rewriting and bug combination. SWE-Mirror (Wang et al., 2025a) further scales this by mirroring issues across repositories. SWE-Flow (Zhang et al., 2025b) utilizes Test-Driven Development (TDD) to produce structured development tasks with explicit goals directly from unit tests, thereby generating training instances.
@@ 3.2 Task Proposal
Once a task is defined, we generate a corresponding detailed checklist that provides a granular explanation and explicitly specifies the requisite unit tests, standard cases, specific assertions, and potential edge cases. This practice offers two primary advantages. First, unit tests serve as a reliable reward signal. Though previous works have explored training software engineering verifier models for inference-time scaling (Pan et al., 2025; Jain et al., 2025b), unit tests remain more trustworthy provided that current language models and agents can generate adequate tests (Chou et al., 2025). Second, the detailed checklist acts as a clear guide, steering the agent in implementing unit tests in the later stage.
@@ 3.5 Functionality Implementation
After the agent completes the implementation, the test suites in the repository are replaced with the original tests produced during the test generation stage, which is a critical step to mitigate the risk of reward hacking, where the agent might otherwise pass tests by directly editing the test code. The original unit tests are subsequently executed to verify the correctness of implementation. This overall process leverages a similar idea to the concept of self-questioning (Chen et al., 2025), as unit test generation and functionality implementation form a mutual verification loop. An issue in either stage, be it a flawed test or an incorrect implementation, will lead to the failure of unit tests, thereby adding to trustworthiness of the generated unit tests. and also robustness of the entire verification process. Notably, we alternate between unit test generation and functionality implementation, forming a loop. This approach ensures that subsequent unit tests are adapted to the the cumulative functionality, which simplifies the implementation the implementation process by removing the need of revisiting or modifying the previously validated code.
@@
Qwen2.5-Coder-7B - 01.81.8   +0.00+0.00 $45.8$   +0.00+0.00 00.720.72   +0.00+0.00 07.557.55   +00.00+00.00 $1.82$   +0.00+0.00
SWE-Gym-7B 491 $10.6$   $\phantom{0}8.8$ $33.8$   12.0-12.0 01.451.45   +0.73+0.73 05.195.19   $\phantom{0}2.36$ $2.51$   +0.69+0.69
R2E-Gym-7B 3.3k $19.0$   +17.2+17.2 00.720.72 +   $0.00$ 02.662.66   $\phantom{0}4.89$ $2.54$   +0.72+0.72
SWE-smith-7B 2.0k $15.2$   +13.4+13.4 00.000.00   0.72-0.72 $12.30$   $\phantom{0}4.75$ $2.62$   +0.80+0.80
SWE-Play-mix-7B 704 $17.0$   +15.2+15.2 06.46.4   39.4-39.4 03.263.26   +2.54+2.54 $30.69$   +23.14+23.14 $2.95$   +1.13+1.13
@@
Qwen2.5-Coder-32B - 07.07.0   +0.00+0.00 09.59.5   +0.00+0.00 09.429.42   +0.00+0.00 $24.81$   +00.00+00.00 $2.31$   +0.00+0.00
SWE-Gym-32B 491 $20.6$   +13.6+13.6 $13.8$   $\phantom{0}4.3$ 03.263.26   6.16-6.16 $10.34$   14.47-14.47 $2.51$   +0.20+0.20
R2E-Gym-32B 3.3k $34.4$   +27.4+27.4 03.263.26   6.16-6.16 $15.04$   $\phantom{0}9.77$ $3.30$   +0.99+0.99
SWE-smith-32B 5.0k $40.2$   +33.2+33.2 $13.77$   +4.35+4.35 $42.15$   +17.34+17.34 $2.62$   +0.31+0.31
SWE-Play-mix-32B 704 $31.2$   +24.2+24.2 06.86.8   $\phantom{0}2.7$ $18.12$   +8.70+8.70 $44.24$   +19.43+19.43 $3.64$   +1.33+1.33
@@ 4 Capability-Specific Adaption
Issue Resolution: SWE-bench (Jimenez et al., 2024) is the canonical benchmark for software engineering tasks, and it features resolving real GitHub issues. This task setting requires a repository that is generally functional but contains a minor bug affecting certain functionalities. With our pipeline, a functional repository already exists, with the only remaining task being the generation of a specific issue and its injection into the code. Similar to Yang et al. (2025b), we first leverage an LLM to formulate an issue based on the unit test documentation and code of the task, and then call an agent to inject the issue into the repository. The existing test suites can directly verify the correctness of the injection, where a failing test serves as the success signal.
Test Generation: SWT-Bench (Mündler et al., 2024) is a test generation benchmark where agents are required to develop unit tests to reproduce issues. For this setting, beyond injecting the issue into the repository, we also require the agent to modify the test suites so that the injected bug is no longer directly detectable. The agents are then tasked with formulating a test script that exposes the faulty behavior and then integrate this script into the existing test suites to complete the task, a procedure identical to SWT-Bench.
Generating Libraries from Scratch: Commit-0 (Zhao et al., 2024) evaluates agents on constructing the entire repository from scratch, with no step-to-step decomposition. For this setting, the complete and functional repositories generated by our pipeline serve as the target artifacts and reference implementation. To simulate the Commit-0 task format, we replace all function bodies with pass statements and then prompt agents to re-implement the project. Notably, we do not use the existing test suites to verify implementation correctness in our experiments. Instead, we let the agent attempt the full task and directly collect its trajectory, which aligns with a pure distillation setting. We adopt such strategy out of the practical necessity. Because our pipeline separates functionalities into step-by-step tasks and the unit tests are designed to be inherently stringent, it is challenging for the agent to pass all tests within a single call.
@@ 5.1 Experimental Settings
With our proposed SWE-Playground pipeline, we collect from 28 generated projects a total of 280 general trajectories (as detailed in Section 3.1 to Section 3.5) and 424 task specific trajectories (as detailed in Section 4), respectively 213 for issue resolution, 183 for issue reproduction and 28 for library generation. Following Pan et al. (2025) and Jain et al. (2025b), we use OpenHands (Wang et al., 2025b) CodeAct Agent our agent Scaffold, and select Qwen-2.5-Coder series (Hui et al., 2024) as our the base model. We adopt SWE-bench Verified (Chowdhury et al., 2024), SWT-Bench Lite and Commit-0 Lite as the evaluation splits for the respective benchmarks. We provide details in Appendix B.
@@
Qwen2.5-Coder 01.81.8 $45.8$ $0.72$ 07.557.55 $1.82$
SWE-Play-mix $17.0$ 06.46.4 $3.26$ $30.69$ $2.95$
SWE-Play-general 08.68.6 $22.6$ $2.54$ 07.557.55 $2.62$
SWE-Play-swe $11.0$ $10.6$ $0.72$ 04.804.80 $2.38$
@@ 6.2 Investigation into Data Efficiency
These outcomes raise a critical question: why can our framework demonstrate better data efficiency compared with various previous methods? To answer this, we perform a detailed comparative analysis of our trajectories against those from baseline studies, with the results presented in Table 4. Our analysis reveals that trajectories generated with SWE-Playground are quantifiably more complex and comprehensive. As shown in the table, they contain two to three times more tokens and messages on average, including a higher number of assistant tokens, tool calls and lines of code edited, compared to other datasets. We characterize this quality as a higher learning density. This concept, in line with the Agency Efficiency Principle introduced by Xiao et al. (2025), suggests that trajectories dense with diverse and complex actions provide a more potent learning impetus per sample. The richness of our trajectories demonstrates this principle in practice, leading to greater data efficiency.